# Priors from Envisioned Posterior Judgments: A Novel Elicitation Approach With Application to Bayesian Clinical Trials

### v2

@@ Abstract
The uptake of formalized prior elicitation from experts in Bayesian clinical trials has been limited, largely due to the challenges associated with complex statistical modeling, the lack of practical tools, and the cognitive burden on experts required to quantify their uncertainty using probabilistic language. Additionally, existing methods do not address prior-posterior coherence, i.e., does the posterior distribution, obtained mathematically from combining the estimated prior with the trial data, reflect the expert’s actual posterior beliefs? We propose a new elicitation approach that seeks to ensure prior-posterior coherence and reduce the expert’s cognitive burden. This is achieved by eliciting responses about the expert’s envisioned posterior judgments under various potential data outcomes and inferring the prior distribution by minimizing the discrepancies between these responses and the expected responses obtained from the posterior distribution. The feasibility and potential value of the new approach are illustrated through an application to a real trial currently underway.
1 Introduction
@@
Bayesian clinical trials have become more popular recently [1, 2, 3]. A key feature of the Bayesian approach is the potential to incorporate domain knowledge, previous data, or expert opinions through prior distributions on model parameters into the design and analysis to increase trial efficiency and obtain more accurate estimates of the intervention effects [4, 5]. However, the adoption of Bayesian methods has been slow despite growing interest [6]. The lack of agreement and methodology on how to choose prior distributions for model parameters is arguably the biggest hurdle to Bayesian analysis [7]. Prior elicitation is a scientific process that transforms domain knowledge, previous data, or expert judgments into well-defined prior distributions. It offers a solution to the prior specification problem, especially when limited data is available [8, 9]. Applied to clinical trials, elicitation involves engaging medical experts to assist them with summarizing their judgments about how well treatments work, based on available data and clinical experience, and communicating those judgments in a way that the results can be combined with trial data [10]. However, a major impediment has been the availability, user acceptability, and difficulty of implementation of approaches to eliciting priors from experts.
@@
Formalized prior elicitation approaches have been available for more than two decades [8]. However, uptake in these methods has been limited in practice due to the need for complex statistical modeling, lack of practical tools and implementation frameworks, and a shortage of trained of elicitation facilitators [11]. Current elicitation approaches include (1) direct elicitation of the probability distribution for the parameter(s), (2) indirect elicitation based on judgments about the plausibility of observable data, and (3) modeling of experts’ past decisions [12, 13, 10, 14]. These methods typically require experts to quantify their uncertainty using probabilistic language. However, while humans typically are comfortable with providing point estimates (i.e., “best guesses”), they are less comfortable with providing probabilistic statements regarding their uncertainty. For example, one may be comfortable stating, “I think there’s a 20% chance it will rain tomorrow.”, but be uneasy about making a statement such as “I am 90% certain the chance of rain tomorrow is between 10% and 30%.”) [15]. Judgments needed for statements like the latter one impose a high cognitive burden on experts [16], which may require them to undergo extensive training to ensure they are adept at quantifying uncertainty using probabilistic statements and to mitigate potential cognitive biases [17].
@@
One previously unacknowledged challenge is that existing methods do not address prior-posterior coherence, i.e., does the posterior distribution, obtained mathematically from combining the elicited prior with the trial data, reflect the expert’s actual posterior beliefs? If one proceeds axiomatically, then on observing new data, the posterior is determined by these data and the elicited prior. However, in a real-world application, it is possible that, based on the data, the expert disagrees with this posterior. If one accepts the axioms and data as correct, then this lack of coherence would indicate that the previously elicited prior did not reflect the expert’s true prior beliefs. A meaningful discrepancy would suggest the elicited prior did not, in fact, reflect the expert’s true prior beliefs. Indeed, one could argue that prior-posterior coherence is the “litmus test” for judging whether the elicited prior truly reflects the expert’s prior beliefs. Yet existing elicitation methods focus solely on obtaining a prior distribution which is never tested in this way.
@@
In this work, we propose a novel elicitation method, referred to as the “prior-from-posteriors (PfP)” approach, that seeks to ensure prior-posterior coherence and reduce the expert’s cognitive burden. This method distinguishes itself from existing approaches in two significant ways: (1) We obtain from experts their envisioned posterior judgments given hypothetical outcome data from the trial, and (2) We collect only point estimates, excluding any judgments regarding certainty (uncertainty). In Section 2, we present our general mathematical framework for inferring the prior distribution, with a focus on elicitation for a univariate, Normally distributed mean parameter. In Section 3, we illustrate the application of our method in a case study - the Walk ’n Watch (WnW) trial, including a description of the implementation aspects. We conclude in Section 4 with a discussion of potential ways to improve the elicitation process and to extend the framework to more general problems.
@@ 2 Methods
  • Let $\mathbf{Y}_{k}$, $k=1,\ldots,K$ denote a set of potential data outcomes (scenarios). For each scenario, we elicit from the expert a point estimate $\tilde{\theta_{k}}$ for the posterior mean $\mu_{k}(a)\equiv E(\theta\mid\mathbf{Y}_{k};a)$.
  • Let $H_{k}(a)\equiv H\left(\tilde{\theta_{k}},\mu_{k}(a)\right)$, where $H(u,v)$ is a chosen discrepancy function; e.g. squared difference $H(u,v)=(u-v)^{2}$. Set the inferred prior distribution to $f(\theta;\hat{a})$ where $\hat{a}$ is the value of $a$ that minimizes the discrepancies on average across all $K$ scenarios, e.g., the root mean square discrepancy RMSD$(a)=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left(\tilde{\theta_{k}}-\mu_{k}(a)\right)^{2}}$.
The value of $H_{k}(\hat{a})$ measures the “incoherency” of the expert’s judgment in scenario $k$ under the inferred prior and can be used by the expert to reflect on and refine their elicitation, and RMSD$(\hat{a})$ reflects the expert’s achieved overall level of (in)coherence (with zero indicating perfect coherence).
In the case study presented below, the goal was to elicit a prior for a (univariate) mean parameter $\theta$ from a Normal data model, i.e., $Y\mid\theta\sim\mathcal{N}\left(\theta,1/\tau\right)$, where $\tau$ is the precision parameter. For the $k$-th scenario, experts were asked to provide their response $\tilde{\theta_{k}}$ given data $\mathbf{Y}_{k}$, comprised of a sample of $n_{k}$ observations for $Y$. We assumed a conjugate Normal prior $\mathcal{N}\left(\mu_{0},1/\tau_{0}\right)$ for $\theta$, and our elicitation goal was to infer the value of $a\equiv(\mu_{0},\tau_{0})$. Under these assumptions, the posterior mean is $\mu_{k}(a)=(\tau_{0}\mu_{0}+n_{k}\bar{y}/s^{2})/(\tau_{0}+n_{k}/s^{2})$, where $\bar{y}$ is the sample mean and $s^{2}$ is the sample variance. With a squared difference as the discrepancy function, the inferred value for ($\mu_{0}$, $\tau_{0}$) from minimizing the RMSD is
@@
3 Case study
3.1 Trial background
Many individuals who have had a stroke experience impaired mobility, and a key therapeutic goal is to regain their ability to walk independently. The Walk ’n Watch (WnW) study (author JE is the trial’s Principal Investigator (PI)) is a stepped-wedge cluster randomized trial (SW-CRT) that aims to investigate the impact of introducing a structured, progressive exercise program called the WnW protocol on the walking ability in adult patients recovering from stroke during their hospital inpatient rehabilitation [18]. The primary outcome in the WnW trial is the 6-Minute Walk Test (6MWT), which measures the distance that a participant can walk in 6 minutes. The trial involved twelve hospital sites, all of whom initially delivered standard care (phase I). Every four months, three hospitals were chosen randomly from those still delivering standard care to switch to delivering the WnW protocol (phase II). A 2-week training period for the WnW protocol was scheduled during the transition from Phase I to Phase II during which enrolment was suspended.
3.2 Motivation for the elicitation
@@ 3.4 Elicitation process
    Circulate introductory elicitation materials/information to the experts
@@
    Host an open online session for the experts to answer their questions about the elicitation
@@
    Collect initial elicitation responses from the experts and analyze to infer their prior distributions
@@
    Collect revised elicitation responses and infer revised prior distributions
@@ 3.4.1 Introductory materials
Scenario Sample Size, $n$ Mean Change in 6MWT, $\bar{y}$ (m)
@@
15 100 $-10$
16 100 $-30$
Table 1: Summary of hypothetical trial outcomes that were presented to experts for elicitation of posterior judgments.
@@ 3.4.3 Initial elicitation data collection and analysis
Experts were asked to complete the questionnaire by providing their best estimate for the mean change in 6WMT from 2021 to 2022 in the UC group (i.e., $\tilde{\theta_{k}}$) if they were to observe the outcomes from each of the 16 hypothetical scenarios. Our statisticians did not directly communicate with the experts during this stage; instead, the PI handled all communication. After receiving the elicited responses, we inferred the prior distribution for each expert and assessed their level of coherency using the method described in section 2.3.
@@ 3.4.4 Feedback to experts
We re-iterate that there is no right or wrong response to any given scenario. However, it is important that your judgments be reasonably consistent across different scenarios. Here are two examples of what we mean by being consistent. Example 1: Suppose that in scenario 1 (where you have not yet seen any outcome data), you think the average change on the 6MWT was $-10$m. In scenario 2, you are shown data that suggests no change. Then, a consistent response from you would be a value between $-10$m and 0m. It would be inconsistent to choose a value less than $-10$m since data showing no change would not support a decrease even larger than what you initially believed. Similarly, it would be inconsistent to choose a value greater than 0m since neither your initial judgment nor the data support a positive change. Example 2: Suppose your responses were 0m (no change) for scenario 1 and $+5$m for scenario 3 (data show +10m based on 10 participants per arm). Then a consistent response for scenario 8 (data show $+10$m based on 30 participants per arm) would be a value in the range $+5$m to $+10$m. The reason is that even with only 10 participants per arm, there was sufficient evidence to shift the expert’s judgment from 0m to 5m, so with stronger evidence (i.e., based on more participants), the shift should be even greater (up to the $+10$m shown in the data). In summary, a simple approach to improving consistency is to ensure your response falls between what you judge the value to be when you have no data and the value shown in the data, with more weight being given to the data as the sample size increases.
@@ 4.1 Initial elicitation
The inferred prior mean and standard deviation (SD) in metres for each expert are displayed in Table 2. This table was ordered from least coherent to most coherent, measured by MSD. There was a large variation in both the means ($-30.8$m to 3.8m) and the SD (0.2m to 118.6m), indicating that beliefs about the time trend were highly diverse. However, the coherency measures indicate that there was poor coherence in the responses for many of the experts, which raises doubt as to whether these priors truly reflect the experts’ beliefs. Figure 1 shows the observed against predicted responses for two experts, one who achieved very good coherence and one who achieved poor coherence.
Figure 2: Comparison of elicited vs. best-fitted responses for experts 2 (left) and expert 9 (right). The labels A to N indicate the responses from scenarios 1 to 16, respectively, and the straight line is the line of equality between the elicited and best-fit response.
Three experts (1, 2, 3) achieved a RMSD less than 3m. Note, though, that the expert who achieved the highest coherency (1) had an inferred prior mean of $-0.3$m and SD of 0.2m, indicating they were nearly certain that the mean change was very close to zero. For this expert, a concern would be whether this degree of certainty could be justified, despite the judgments being coherent. Examination of the individual responses for expert 10 strongly suggested the expert misunderstood what was requested, which would explain their large values for the prior SD and the incoherency measures.
@@
Mean, m SD, m RMSD, m
Expert 1 $-0.3$ 0.2 1.2
Expert 2 $-0.3$ 10.4 1.3
Expert 3 $-30.8$ 4.3 1.7
Expert 4 $-9.4$ 11.6 3.4
@@
Expert 6 0 32.2 6.0
Expert 7 0 15.2 6.1
Expert 8 $-2.6$ 30.6 6.6
Expert 9 $-15.8$ 16 8.7
@@
Table 2: The mean, SD, and coherency (RMSD) of inferred Normal prior distribution for each expert.
@@ 4.2 Revised elicitation
Inferred Mean, m Inferred SD, m RMSD, m
Initial Revised Initial Revised Initial Revised
Expert 1 $-0.3$ 0 0.2 0 1.2 0.0
Expert 2* $-0.3$ $-0.3$ 10.4 10.4 1.3 1.3
Expert 3 $-30.8$ $-30.9$ 4.3 5.1 1.7 1.1
Expert 4* $-9.4$ $-9.4$ 11.6 11.6 3.4 3.4
Expert 5* 3.8 3.8 20.9 20.9 5.5 5.5
Expert 6 0 0 32.2 40.9 6.0 3.9
Expert 7 0 0 15.2 7.4 6.1 1.6
Expert 8 $-2.6$ $-3.3$ 30.6 32.1 6.7 6.3
Expert 9 $-15.8$ $-12$ 16 15.5 8.7 2.8
Expert 10 0.9 $-26.7$ 118.6 21.3 21.7 3.7
Table 3: Comparison of Initial and Revised means, SDs, and coherencies (RMSDs) of the inferred Normal prior distributions for each expert
* Denotes expert who did not submit revised responses

### v3

@@ Abstract
Background: The uptake of formalized prior elicitation from experts in Bayesian clinical trials has been limited due to challenges such as complex statistical modeling, lack of practical tools, and the cognitive burden placed on experts arising from needing to quantify their uncertainty probabilistically. Existing methods also fail to address prior-posterior coherence, i.e., how do we ensure that the posterior distribution, obtained mathematically from combining the estimated prior with the trial data, reflects the expert’s actual posterior beliefs?
Method: In this study, we propose a new elicitation approach that effectuates prior-posterior coherence and reduces cognitive burden. This is achieved by eliciting expert responses, comprising point estimates only, about envisioned posterior judgments under various data outcomes and inferring the prior distribution by minimizing discrepancies between these responses and expected responses derived from the posterior distribution. Via an iterative process, experts receive feedback on the degree of coherency of their responses, and are invited to revise their responses to achieve greater coherency. The feasibility and potential value of this new approach are illustrated through an application to an ongoing trial.
Results: We involved 10 experts from Walk ’n watch trial research team. Experts were presented with 16 hypothetical outcome scenarios to experts and elicit the priors followed by the developed elicitation framework. Following two rounds of elicitation, experts’ judgments showed substantial improvement in coherency, demonstrating the practical applicability of the proposed elicitation approach.
Conclusion: The proposed method provides a practical solution to the challenges of formalized prior elicitation in Bayesian clinical trials by addressing prior-posterior coherence and reducing cognitive demands on experts.
Keywords: Prior elicitation, Bayesian, Decision making, Trial design, Informative prior, Prior-posterior coherence
1 Background
@@
Bayesian clinical trials have become more popular recently [1, 2, 3]. A key feature of the Bayesian approach is the potential to incorporate domain knowledge, previous data, or expert opinions through prior distributions on model parameters into the design and analysis to increase trial efficiency and obtain more accurate estimates of the intervention effects [4, 5]. However, the adoption of Bayesian methods has been slow despite growing interest [6]. The lack of agreement and methodology on how to choose prior distributions for model parameters is arguably the biggest hurdle to Bayesian analysis [7]. Prior elicitation is a scientific process that transforms domain knowledge, previous data, or expert judgments into well-defined prior distributions. It offers a solution to the prior specification problem, especially when limited data is available [8, 9]. Applied to clinical trials, elicitation involves engaging medical experts to assist them in summarizing their judgments on how well treatments work, based on available data and clinical experience, and communicating those judgments in a way that the results can be combined with trial data [10]. However, a major impediment has been the availability, user acceptability, and difficulty of implementation of approaches to eliciting priors from experts.
@@
Formalized prior elicitation approaches have been available for more than two decades [8]. However, uptake in these methods has been limited in practice due to the need for complex statistical modeling, lack of practical tools and implementation frameworks, and a shortage of trained of elicitation facilitators [11]. Current elicitation approaches include (1) direct elicitation of the probability distribution for the parameter(s), (2) indirect elicitation based on judgments about the plausibility of observable data, and (3) modeling of past experts’ decisions [12, 13, 10, 14]. These methods typically require experts to quantify their uncertainty using probabilistic language. However, while humans typically are comfortable with providing point estimates (i.e., “best guesses”), they are less comfortable with providing probabilistic statements regarding their uncertainty. For example, one may be comfortable stating “I think there is a 20% chance that it will rain tomorrow.”, but be uneasy about making a statement such as “I am 90% certain that the chance of rain tomorrow is between 10% and 30%.”) [15]. The judgments needed for statements like the latter impose a high cognitive burden on experts [16], which may require them to undergo extensive training to ensure they are adept at quantifying uncertainty using probabilistic statements and to mitigate potential cognitive biases [17].
@@
One previously unacknowledged challenge is that existing methods do not address prior-posterior coherence, i.e., does the posterior distribution, obtained mathematically from combining the elicited prior with the trial data, reflect the expert’s actual posterior beliefs? If one proceeds axiomatically, then on observing the trial data, the posterior calculated from these data and the elicited prior describes what the expert ought to believe. However, it is possible that after seeing the trial data, the expert may decide that this “mathematical” posterior does not describe the expert’s actual posterior beliefs. If one accepts the axioms and data as correct (and assuming the expert had no reason to modify their prior), then a meaningful discrepancy between these posteriors would suggest that the previously elicited prior did not, in fact, reflect the true prior beliefs of the expert. Indeed, one could argue that prior-posterior coherence is the “litmus test” for judging whether the elicited prior truly reflects the expert’s prior beliefs. Yet existing elicitation methods focus solely on obtaining a prior distribution which is never tested in this way.
@@
In this work, we propose a novel elicitation method, referred to as the “prior-from-posteriors (PfP)” approach, which seeks to ensure prior-posterior coherence and reduce the cognitive burden of the expert. This method distinguishes itself from existing approaches in two significant ways: (1) We obtain from experts their envisioned posterior judgments given hypothetical outcome data from the trial, and (2) We collect only point estimates, excluding any judgments regarding certainty (uncertainty). In Section 2, we present our general mathematical framework for inferring the prior distribution, with a focus on elicitation for a univariate, normally distributed mean parameter. In Section 3, we illustrate the application of our method in a case study - the Walk ’n Watch (WnW) trial, including a description of the implementation aspects. In Section 4, we conclude with a discussion of potential ways to improve the elicitation process and to extend the framework to more general problems.
@@ 2 Methods
  • Let $\mathbf{Y}_{k}$, $k=1,\ldots,K$ denote a set of potential data outcomes (scenarios). For each scenario, we elicit from the expert a point estimate $\tilde{\theta_{k}}$ for the posterior mean
    $\mu_{k}(a)\equiv E(\theta\mid\mathbf{Y}_{k};a).$
[…]
    $RMSD(a)=\sqrt{\frac{1}{K}\sum_{k=1}^{K}H\left(\tilde{\theta_{k}},\mu_{k}(a)\right)}=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left(\tilde{\theta_{k}}-\mu_{k}(a)\right)^{2}}.$
[…]
In the case study presented below, the objective was to obtain a prior for the mean parameter (univariate) $\theta$ from a normal data model, i.e., $Y\mid\theta\sim\mathcal{N}\left(\theta,1/\tau\right)$, where $\tau$ is the precision parameter. For the $k$-th scenario, experts were asked to provide their response $\tilde{\theta_{k}}$ given data $\mathbf{Y}_{k}$, comprised of a sample of $n_{k}$ observations for $Y$. We assumed a conjugate normal prior $\mathcal{N}\left(\mu_{0},1/\tau_{0}\right)$ for $\theta$, and our elicitation goal was to infer the value of $a\equiv(\mu_{0},\tau_{0})$. Under these assumptions, the posterior mean is $\mu_{k}(a)=(\tau_{0}\mu_{0}+n_{k}\bar{y}/s^{2})/(\tau_{0}+n_{k}/s^{2})$, where $\bar{y}$ is the sample mean and $s^{2}$ is the sample variance. With a squared difference as the discrepancy function, the inferred value for ($\mu_{0}$, $\tau_{0}$) from minimizing the RMSD is
@@
2.1 Case Study
2.1.1 Trial background
Many individuals who have had a stroke experience impaired mobility, and a key therapeutic goal is to regain their ability to walk independently. The Walk ’n Watch (WnW) study (author JE is the trial’s Principal Investigator (PI)) is a stepped-wedge cluster randomized trial (SW-CRT) that aims to investigate the impact of introducing a structured, progressive exercise program called the WnW protocol on the walking ability in adult patients recovering from stroke during their hospital inpatient rehabilitation [18, 19]. The primary outcome in the WnW trial is the 6-Minute Walk Test (6MWT), which measures the distance that a participant can walk in 6 minutes. The trial involved twelve hospital sites, all of whom initially delivered standard care (phase I). Every four months, three hospitals were chosen randomly from those still delivering standard care to switch to delivering the WnW protocol (Phase II). A 2-week training period for the WnW protocol was scheduled during the transition from Phase I to Phase II, during which enrollment was suspended.
2.1.2 Motivation for the elicitation
@@ 2.2 Elicitation process
    Circulate introductory elicitation materials/information to the experts.
@@
    Host an open online session for the experts to answer their questions about the elicitation.
@@
    Collect initial elicitation responses from the experts and analyze to infer their prior distributions.
@@
    Collect revised elicitation responses and infer revised prior distributions.
@@ 2.2.1 Introductory materials
Table 1: Summary of hypothetical trial outcomes that were presented to experts for elicitation of posterior judgments
@@
Scenario Sample Size Mean 6MWT Change
($n$) ($\bar{y}$, meters)
@@ 1 No Data Not Applicable
15 100 10-10
16 100 30-30
@@ 2.2.3 Initial elicitation data collection and analysis
Experts were asked to complete the questionnaire by providing their best estimate for the mean change in 6WMT from 2021 to 2022 in the UC group (i.e., $\tilde{\theta_{k}}$) if they were to observe the outcomes from each of the 16 hypothetical scenarios. Our statisticians did not directly communicate with the experts during this stage; instead, the PI handled all communication. After receiving the elicited responses, we inferred the prior distribution for each expert and assessed their level of coherency using the method described in section 2.
@@ 2.2.4 Feedback to experts
We re-iterate that there is no right or wrong response to any given scenario. However, it is important that your judgments be reasonably consistent across different scenarios. Here are two examples of what we mean by being consistent. Example 1: Suppose that in scenario 1 (where you have not yet seen any outcome data), you think the average change on the 6MWT was 10-10m. In scenario 2, you are shown data that suggests no change. Then, a consistent response from you would be a value between 10-10m and 0m. It would be inconsistent to choose a value less than 10-10m since data showing no change would not support a decrease even larger than what you initially believed. Similarly, it would be inconsistent to choose a value greater than 0m since neither your initial judgment nor the data support a positive change. Example 2: Suppose your responses were 0m (no change) for scenario 1 and +5+5m for scenario 3 (data show +10m based on 10 participants per arm). Then a consistent response for scenario 8 (data show +10+10m based on 30 participants per arm) would be a value in the range +5+5m to +10+10m. The reason is that even with only 10 participants per arm, there was sufficient evidence to shift the expert’s judgment from 0m to 5m, so with stronger evidence (i.e., based on more participants), the shift should be even greater (up to the +10+10m shown in the data). In summary, a simple approach to improving consistency is to ensure your response falls between what you judge the value to be when you have no data and the value shown in the data, with more weight being given to the data as the sample size increases.
@@ 3.1 Initial elicitation
The inferred prior mean and standard deviation (SD) in metres for each expert are displayed in Table 2. This table was ordered from least coherent to most coherent, measured by RMSD. There was a large variation in both the means (30.8-30.8m to 3.8m) and the SD (0.2m to 118.6m), indicating that beliefs about the time trend were highly diverse. However, the coherency measures indicate that there was poor coherence in the responses for many of the experts, which raises doubt as to whether these priors truly reflect the experts’ beliefs. Figure 1 shows the observed versus predicted responses for two experts, one who achieved very good coherence and one who achieved poor coherence.
Figure 2: Comparison of elicited vs. best-fitted responses for experts 2 (left:2A) and expert 9 (right:2B). The labels A to N indicate the responses from scenarios 1 to 16, respectively, and the straight line is the line of equality between the elicited and best-fit response.
Three experts (1, 2, 3) achieved a RMSD less than 3m. Note, though, that the expert who achieved the highest coherency (1) had an inferred prior mean of 0.3-0.3m and SD of 0.2m, indicating they were nearly certain that the mean change was very close to zero. For this expert, a concern would be whether this degree of certainty could be justified, despite the judgments being coherent. Examination of the individual responses for expert 10 strongly suggested the expert misunderstood what was requested, which would explain their large values for the prior SD and for the incoherency measure.
Table 2: The mean, SD, and coherency (RMSD) of the inferred normal prior distribution for each expert (in meters).
@@
Mean SD RMSD
Expert 1 0.3-0.3 0.2 1.2
Expert 2 0.3-0.3 10.4 1.3
Expert 3 30.8-30.8 4.3 1.7
Expert 4 9.4-9.4 11.6 3.4
@@
Expert 6 0.0 32.2 6.0
Expert 7 0.0 15.2 6.1
Expert 8 2.6-2.6 30.6 6.6
Expert 9 15.8-15.8 16.0 8.7
@@ 3.2 Revised elicitation
[b] Inferred Mean Inferred SD RMSD Initial Revised Initial Revised Initial Revised Expert 1 0.3-0.3 0.0 0.2 0.0 1.2 0.0 Expert 2* 0.3-0.3 0.3-0.3 10.4 10.4 1.3 1.3 Expert 3 30.8-30.8 30.9-30.9 4.3 5.1 1.7 1.1 Expert 4* 9.4-9.4 9.4-9.4 11.6 11.6 3.4 3.4 Expert 5* 3.8 3.8 20.9 20.9 5.5 5.5 Expert 6 0.0 0.0 32.2 40.9 6.0 3.9 Expert 7 0.0 0.0 15.2 7.4 6.1 1.6 Expert 8 2.6-2.6 3.3-3.3 30.6 32.1 6.7 6.3 Expert 9 15.8-15.8 12-12 16.0 15.5 8.7 2.8 Expert 10 0.9 26.7-26.7 118.6 21.3 21.7 3.7
Table 3: Comparison of Initial and Revised means, SDs, and coherencies (RMSDs) of the inferred normal prior distributions for each expert (in meters)
    Expert who did not submit revised responses